# A Versatility Measure for Parametric Risk Models

### v1

@@ 5.1 Maximum-Entropy Prior
In the present context, all we know about the parameter vector $\boldsymbol{a}$ is that it should be representative of all possible vectors employed to characterize either a frequency or severity distribution. Since most frequency PMFs and severity PDFs are commonly formulated (or easily reformulated) with real-valued parameters spanning the space $\mathbb{R}^{+}$ (as in the Exponential case, where both the mean, $\theta$, and its inverse, $\lambda$, can be any positive real number), we will restrict attention to vectors $\boldsymbol{a}$ spanning $\left(\mathbb{R}^{+}\right)^{k}$ and seek a reasonable uninformative prior distribution according to the principle of indifference. This usually takes one of two approaches: the method of Jeffreys priors, or that of maximum-entropy priors. [^16] See Jaynes (2003) for an extensive discussion of uninformative prior distributions.
@@
$\left[b_{1},\ldots,b_{k}\right]=\boldsymbol{\textrm{g}}\left(a_{1},\ldots,a_{k}\right)$
@@
$p_{\textrm{J}}^{\prime}\left(b_{1},\ldots,b_{k}\right)=p_{\textrm{J}}\left(a_{1},\ldots,a_{k}\right)\left|\det\left(\boldsymbol{\textrm{J}}\right)\right|,$
@@
To overcome this problem, we can require the distribution of $\ln\left(a\right)\in\mathbb{R}$ to be symmetric about 0. Here, the maximum-entropy distribution, subject to a fixed variance, $\sigma^{2}$, is $\textrm{Lognormal}\left(\mu=0,\sigma\right)$, with $\sigma\in\mathbb{R}^{+}$, and we can set the hyper-parameter $\sigma$ equal to 1 based on the rationale that $\ln\left(\sigma\right)$, like $\ln\left(a\right)$, must be symmetric about the origin (which, for a constant, is exactly 0). Taking the parameters $a_{1},\ldots,a_{k}$ in (18) to be IID $\textrm{Lognormal}\left(\mu=0,\sigma=1\right)$ random variables with joint distribution function $P_{\textrm{LN}}\equiv P_{\textrm{LN}}\left(\boldsymbol{a}\right)$ then yields the following expression:
@@
Figure 1. PDFs of $\textrm{Exponential}\left(\lambda\right)$, $\textrm{Gamma}\left(r,\lambda=1\right)$, and $\textrm{Weibull}\left(\lambda=1,\tau\right)$ Distributions
(parameter $=0.5$ for dotted line, $1.0$ for solid line, and $2.0$ for dashed line)
@@ 5.2 Least-Compressible Parameterization
$\bullet$ Other special functions, as well as more complicated mathematical operations (e.g., summations, limits, derivatives, integrals, etc.) are not permitted.
@@ 6 Illustrative Applications
Having defined a formal measure of functional versatility, we now apply this measure to eight distinct families of 2-parameter distributions to explore its implications. Of these eight families, four are defined on $\mathbb{R}_{0}^{+}$ (i.e., appropriate for severities) and four on $\mathbb{Z}_{0}^{+}$ (i.e., appropriate for frequencies).
@@ 6.1 Continuous Distributions
The four continuous families – $\textrm{Gamma}\left(r,\lambda\right)$, $\textrm{Weibull}\left(\lambda,\tau\right)$, Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$, and $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma\right)$ – are selected primarily for their common use as severity models. Their PDFs and versatility measures are presented in Table 3, with all parameters taken to span $\mathbb{R}^{+}$. To provide some (admittedly limited) sense of the impact of individual parameters on the versatility measure, the table includes two 1-parameter special cases of each 2-parameter family (three of which are identical to the 1-parameter examples shown in Table 2 of Subsection 5.1).
@@
  • For three of the four families, the versatility measure of the 2-parameter distribution falls between the measures of the two 1-parameter special cases. This property is somewhat intuitive because one would expect the distributional characteristics contributed by each of the two parameters individually to be “averaged” when taken together. Therefore, it is important to note that the property does not hold for all distributions. For $\textrm{Gamma}\left(r,\lambda\right)$, the fact that
    $\mathcal{V}\left(F_{Y\mid r,\lambda}^{\left(\Gamma\right)}\right)>\max\left\{\mathcal{V}\left(F_{Y\mid r=1,\lambda}^{\left(\Gamma\right)}\right),\mathcal{V}\left(F_{Y\mid r,\lambda=1}^{\left(\Gamma\right)}\right)\right\}$
    suggests that the $r$ and $\lambda$ parameters positively reinforce each other’s impact on versatility. However, the analysis of such marginal contributions to versatility lies beyond the scope of the present work.
@@ 6.2 Discrete Distributions
(3) $\textrm{Waring}\left(\alpha,\vartheta\right)$ is characterized by strictly decreasing PMFs with heavy tails (such that for all real $\kappa\leq\alpha$, $\textrm{E}_{X\mid\alpha,\vartheta}\left[X^{\kappa}\right]=\infty$), just as Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ is characterized by strictly decreasing PDFs with heavy tails (such that for all real $\kappa\leq\alpha$, $\textrm{E}_{Y\mid\alpha,\vartheta}\left[Y^{\kappa}\right]=\infty$).
@@
A comparison of Tables 3 and 4 reveals:
  • For the Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}\right)$ family with $\textrm{E}_{X\mid r,\tfrac{m}{m+1}}\left[X\right]=\tfrac{1}{m}$, there is a second, equally compressible parameterization, Negative $\textrm{Negative Binomial}\left(r,\tfrac{1}{m+1}\right)$, for which $\textrm{E}_{X\mid r,\tfrac{1}{m+1}}\left[X\right]=m$. This is analogous to the fact that $\textrm{Gamma}\left(r,\lambda\right)$ possesses the equally compressible counterpart $\textrm{Gamma}\left(r,\theta\right)$ in the continuous case (as noted in Subsection 5.2). However, unlike the continuous case, the two different parameterizations yield different versatility measures, which we resolve by averaging (as suggested in Subsection 5.2). The same issue arises for the Discrete $\textrm{Discrete Weibull}\left(\tfrac{m}{m+1},\tau\right)$ family, and is handled in the same way. [^20] The versatility measures of distributions parameterized by $\tfrac{1}{m+1}$ are noticeably smaller than those of distributions based on $\tfrac{m}{m+1}$ because the former PMFs are particularly sensitive to $m$ for values of $x$ close to 00, whereas the latter are relatively more sensitive as $x\rightarrow\infty$. For the discrete sample space $x\in\mathbb{Z}_{0}^{+}$, there is little opportunity for a PMF to benefit from functional adaptability in the neighborhood of 00, hence less functional versatility.
@@
  • The versatility measures associated with the first three discrete families are consistently lower than the corresponding measures of their continuous analogues. This arises from the same phenomenon mentioned in Footnote 20: the discrete sample space $x\in\mathbb{Z}_{0}^{+}$ affords dramatically less opportunity for functional adaptability in the neighborhood of 00 than does the continuous sample space $y\in\mathbb{R}_{0}^{+}$.
  • For all four families, the versatility measure of the 2-parameter distribution falls between the measures of the two 1-parameter special cases.
  • The parameters with a direct effect on tail behavior – that is, $\tau$ for the averaged Discrete $\textrm{Weibull}\left(\tfrac{m}{m+1},\tau\right)$ and Discrete $\textrm{Discrete Weibull}\left(\tfrac{1}{m+1},\tau\right)$ distributions and $\alpha$ for the $\textrm{Waring}\left(\alpha,\vartheta\right)$ distribution – appear to have relatively larger effects on versatility than other parameters.
@@
  • The versatility measures of Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$ and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$ are considerably lower than other measures in the table. This is because the two PMFs are quite flat for all values of $r$ and $\vartheta$, respectively, with both functions’ simplicity offset by limited adaptability. As noted in conjunction with $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$ in the previous subsection, we do not propose a specific minimum cutoff for the versatility measure; however, it appears that both $\mathcal{V}\left(F_{X\mid r,\tfrac{m}{m+1}=\tfrac{1}{2}}^{\left(\textrm{NB}\right)}\right)=1.0151$ and $\mathcal{V}\left(F_{X\mid\alpha=1,\vartheta}^{\left(\textrm{War}\right)}\right)=0.9423$ are strong indicators of idiosyncrasy.

### v3

@@ 5.1 Maximum-Entropy Prior
In the present context, all we know about the parameter vector $\boldsymbol{a}$ is that it should be representative of all possible vectors employed to characterize either a frequency or severity distribution. Since most frequency PMFs and severity PDFs are commonly formulated (or easily reformulated) with real-valued parameters spanning the space $\mathbb{R}^{+}$ (as in the Exponential case, where both the mean, $\theta$, and its inverse, $\lambda$, can be any positive real number), we will restrict attention to vectors $\boldsymbol{a}$ spanning $\left(\mathbb{R}^{+}\right)^{k}$ and seek a reasonable uninformative prior distribution according to the principle of indifference. This usually takes one of two approaches: the method of Jeffreys priors, or that of maximum-entropy priors. [^17] See Jaynes (2003) for a classic treatment of uninformative prior distributions and Llorente et al. (2023) for some contemporary perspectives.
@@
$\left[b_{1},b_{2},\ldots,b_{k}\right]=\boldsymbol{\textrm{g}}\left(a_{1},a_{2},\ldots,a_{k}\right)$
@@
$p_{\textrm{J}}^{\prime}\left(b_{1},b_{2},\ldots,b_{k}\right)=p_{\textrm{J}}\left(a_{1},a_{2},\ldots,a_{k}\right)\left|\det\left(\boldsymbol{\textrm{J}}\right)\right|,$
@@
To overcome this problem, we can require the distribution of $\ln\left(a\right)\in\mathbb{R}$ to be symmetric about 0. Here, the maximum-entropy distribution, subject to a fixed variance, $\sigma^{2}$, is $\textrm{Lognormal}\left(\mu=0,\sigma\right)$, with $\sigma\in\mathbb{R}^{+}$, and we can set the hyper-parameter $\sigma$ equal to 1 based on the rationale that $\ln\left(\sigma\right)$, like $\ln\left(a\right)$, must be symmetric about the origin (which, for a constant, is exactly 0). Taking the parameters $a_{1},a_{2},\ldots,a_{k}$ in (18) to be IID $\textrm{Lognormal}\left(\mu=0,\sigma=1\right)$ random variables with joint distribution function $P_{\textrm{LN}}\equiv P_{\textrm{LN}}\left(\boldsymbol{a}\right)$ then yields the following expression:
@@
Since $\textrm{E}_{\textrm{LN}}\left[\left|a_{1}^{p_{1}}a_{2}^{p_{2}}\cdots a_{k}^{p_{k}}\right|\right]<\infty$ for all $\boldsymbol{p}\in\mathbb{R}^{k}$, the right-hand side of (19) is well defined as long as the individual elements of $\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)$ do not grow faster than $\left|a_{1}^{p_{1}}a_{2}^{p_{2}}\cdots a_{k}^{p_{k}}\right|$ as the individual $a_{i}$ approach 00 or $\infty$.
@@
Figure 1. PDFs of the $\textrm{Exponential}\left(\lambda\right)$, $\textrm{Gamma}\left(r,\lambda=1\right)$, and $\textrm{Weibull}\left(\lambda=1,\tau\right)$ Distributions (from left to right)
[variable parameter $=0.5$ for dotted line, $1.0$ for solid line, and $2.0$ for dashed line]
Given the definition in (19), it is clear that $\mathcal{V}\left(F_{0}\right)\in\mathbb{R}^{+},$ and thus not bounded above. This makes the interpretation of versatility magnitudes somewhat challenging, and raises the natural question of whether it is possible to transpose $\mathcal{V}\left(F_{0}\right)$ to a bounded interval, such as $\left(0,1\right)$, in a way that affords meaningful interpretations (as in the case of the regression $R^{2}$, which can be viewed as the proportion of variation explained by a posited regression model). One salient approach would be to consider transformations $\textrm{T}\left(\mathcal{V}\left(F_{0}\right)\right):\mathbb{R}^{+}\rightarrow\left(0,1\right)$ that endow a percentile to each possible versatility measure. In other words, if we could define and identify a theoretical distribution, $F_{\mathcal{V}}\left(v\right)$, of $\mathcal{V}\left(F_{0}\right)$ over all probability distributions $F_{0}$ within some prescribed category (say, e.g., severity distributions on $\mathbb{R}_{0}^{+}$), then we might set $\textrm{T}\left(\mathcal{V}\left(F_{0}\right)\right)=100F_{\mathcal{V}}^{-1}\left(\mathcal{V}\left(F_{0}\right)\right)$.
Although the problem of identifying a suitable theoretical distribution $F_{\mathcal{V}}\left(v\right)$ appears largely intractable, there are potential alternatives. For example, one could employ an uninformative prior on $\mathbb{R}^{+}$ (such as $\textrm{Lognormal}\left(\mu=0,\sigma=1\right)$), or base the choice of $F_{\mathcal{V}}\left(v\right)$ on empirical, rather than theoretical, considerations. In fact, the latter approach is likely to arise naturally in applied work by observing a large number of versatility measures for a certain category of distributions $F_{0}$, yielding an empirical distribution of $\mathcal{V}\left(F_{0}\right)$.
For purposes of the current study, we will defer a deeper investigation of this issue to future research, recognizing that actual computations of $\mathcal{V}\left(F_{0}\right)$ are useful primarily for comparing the characteristics of different probability distributions, rather than assessing the qualities of any single distribution in isolation. In this way, the proposed versatility measure is comparable to other unbounded assessment measures, such as Akaike’s Information Criterion and Mallows’ $C_{p}$ Statistic.
@@ 5.2 Least-Compressible Parameterization
$\bullet$ Other special functions, as well as more complicated mathematical operations (e.g., summations, limits, derivatives, integrals, etc.) are not permitted. [^20] We note that this rule precludes applications to mixture distributions, which are commonly used in risk analysis. However, two practical considerations mitigate the impact of this restriction: (1) it remains feasible to compute and compare the versatility measures of different mixture distributions whose formulations in terms of integrals/sums over the mixing distribution are carried out in a consistent manner; and (2) it is possible to assess the versatility of a mixture distribution informally by measuring the versatilities of its individual components (i.e., the kernel and mixing distribution) separately.
@@ 6 Illustrative Applications
Having defined a formal measure of functional versatility, we now apply this measure to eight distinct families of 2-parameter distributions to explore its implications. Four of these families are defined on $\mathbb{R}_{0}^{+}$, and thus appropriate for modeling severities, and the other four are defined on $\mathbb{Z}_{0}^{+}$, and appropriate for frequencies. Following this overview, we will investigate one of the continuous families more closely to show how the versatility measure can be employed in practice to select certain probability distributions over others.
@@ 6.1 Continuous Distributions
The four continuous families – $\textrm{Gamma}\left(r,\lambda\right)$, $\textrm{Weibull}\left(\lambda,\tau\right)$, Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$, and $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma\right)$ – are chosen primarily for their common use as severity models. Their PDFs and versatility measures are presented in Table 3, with all parameters taken to span $\mathbb{R}^{+}$. To provide some (admittedly limited) sense of the impact of individual parameters on the versatility measure, the table includes two 1-parameter special cases of each 2-parameter family (three of which are identical to the 1-parameter examples shown in Table 2 of Subsection 5.1).
@@
  • For three of the four families, the versatility measure of the 2-parameter distribution falls between the measures of the two 1-parameter special cases. This property seems intuitive because one would expect the distributional characteristics contributed by each of the two parameters individually to be “averaged” when taken together. However, it is important to note that this property does not always hold for extreme values of the “fixed” parameter in the 1-parameter distributions. For example, as the fixed parameter $\alpha^{*}\rightarrow\infty$ in the 1-parameter Pareto $\textrm{Pareto 2}\left(\alpha^{*},\vartheta\right)$, the versatility measure converges to $2.7183$ (which is greater than the measure associated with Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$). Furthermore, as the fixed parameter $\sigma^{*}\rightarrow 0$ in the 1-parameter $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma^{*}\right)$, the versatility measure diverges to $\infty$ (well above the measure associated with $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma\right)$). Apart from these cautionary observations, the analysis of marginal contributions to versatility from individual parameters lies beyond the scope of the present work.
@@ 6.2 Discrete Distributions
(3) $\textrm{Waring}\left(\alpha,\vartheta\right)$ is characterized by strictly decreasing PMFs with heavy tails (such that for all real $\kappa\leq\alpha$, $\textrm{E}_{X\mid\alpha,\vartheta}^{\left(\textrm{War}\right)}\left[X^{\kappa}\right]=\infty$), just as Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ is characterized by strictly decreasing PDFs with heavy tails (such that for all real $\kappa\leq\alpha$, $\textrm{E}_{Y\mid\alpha,\vartheta}^{\left(\textrm{P2}\right)}\left[Y^{\kappa}\right]=\infty$).
@@
From Table 4, we make the following observations:
  • For the Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}\right)$ family with $\textrm{E}_{X\mid r,\tfrac{m}{m+1}}\left[X\right]=\tfrac{1}{m}$, there is a second, equally compressible parameterization, Negative $\textrm{Negative Binomial}\left(r,\tfrac{1}{m+1}\right)$, for which $\textrm{E}_{X\mid r,\tfrac{1}{m+1}}\left[X\right]=m$. This is analogous to the fact that $\textrm{Gamma}\left(r,\lambda\right)$ possesses the equally compressible counterpart $\textrm{Gamma}\left(r,\theta\right)$ in the continuous case (as noted in Subsection 5.2). However, unlike the continuous case, the two different parameterizations yield different versatility measures, which we resolve by averaging (as suggested in Subsection 5.2). The same issue arises for the Discrete $\textrm{Discrete Weibull}\left(\tfrac{m}{m+1},\tau\right)$ family, and is handled in the same way. The versatility measures associated with the first three discrete families are consistently lower than the corresponding measures of their continuous analogues. This arises from the same phenomenon mentioned in Footnote 20: the discrete sample space $x\in\mathbb{Z}_{0}^{+}$ affords dramatically less opportunity for functional adaptability in the neighborhood of 00 than does the continuous sample space $y\in\mathbb{R}_{0}^{+}$.
  • For all four families, the versatility measure of the 2-parameter distribution falls between the measures of the two 1-parameter special cases. However, as noted with regard to Table 3, this property does not always hold for extreme values of the fixed parameter in the 1-parameter distributions.
  • The parameters with a direct effect on tail behavior – that is, $\tau$ for the averaged Discrete $\textrm{Weibull}\left(\tfrac{m}{m+1},\tau\right)$ and Discrete $\textrm{Discrete Weibull}\left(\tfrac{1}{m+1},\tau\right)$ distributions and $\alpha$ for the $\textrm{Waring}\left(\alpha,\vartheta\right)$ distribution – appear to have relatively larger effects on versatility than other parameters. [^22] The versatility measures of distributions parameterized by $\tfrac{1}{m+1}$ are noticeably smaller than those of distributions based on $\tfrac{m}{m+1}$ because the former PMFs are particularly sensitive to $m$ for values of $x$ close to 00, whereas the latter are relatively more sensitive as $x\rightarrow\infty$. For the discrete sample space $x\in\mathbb{Z}_{0}^{+}$, there is little opportunity for a PMF to benefit from functional adaptability in the neighborhood of 00, hence less functional versatility.
@@
  • The versatility measures of Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$ and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$ are noticeably lower than other measures in the table. This is because the two PMFs are quite flat for all values of $r$ and $\vartheta$, respectively, with both functions’ simplicity offset by limited adaptability. As noted in conjunction with $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$ in the previous subsection, we do not propose a specific minimum cutoff for the versatility measure; however, it appears that both $\mathcal{V}\left(F_{X\mid r,\tfrac{m}{m+1}=\tfrac{1}{2}}^{\left(\textrm{NB}\right)}\right)=1.0151$ and $\mathcal{V}\left(F_{X\mid\alpha=1,\vartheta}^{\left(\textrm{War}\right)}\right)=0.9423$ are strong indicators of idiosyncrasy.
6.3 Pareto 2$\left(\alpha,\vartheta\right)$ vs. Inverse Gamma$\left(r,\lambda\right)$
In Subsection 6.1, we noted that the 2-parameter Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ family possesses a rather low versatility measure compared to the other 2-parameter continuous distributions in Table 3. This observation affords the opportunity to consider whether, despite its common use, Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ is unduly idiosyncratic. Given the distribution’s obvious simplicity, the basic question is whether it is insufficiently adaptable for modeling severities effectively, and therefore should be replaced by other distributions with similar characteristics (in particular, heavy tails) but greater adaptability.
One potential alternative is the Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ distribution, associated with the random variable $Y$ such that $\dfrac{1}{Y}\sim\textrm{Gamma}\left(r,\lambda\right)$. The Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ PDF is given by
$f_{Y\mid r,\lambda}^{\left(\textrm{I}\Gamma\right)}\left(y\right)=\dfrac{\lambda^{r}e^{-\lambda/y}}{y^{r+1}\Gamma\left(r\right)},$
where the tail parameter, $r$, plays a role comparable to the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$’s $\alpha$ in that for all real $\kappa\leq r$, $\textrm{E}_{Y\mid r,\lambda}^{\left(\textrm{I}\Gamma\right)}\left[Y^{\kappa}\right]=\infty$. As illustrated in Figure 2, the PDFs of the two distributions can differ considerably near the lower end of the sample space, with Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ possessing a mode at 0 (where $f_{Y\mid\alpha,\vartheta}^{\left(\textrm{P2}\right)}\left(0\right)=\tfrac{\alpha}{\vartheta}>0$) and Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ increasing from $f_{Y\mid r,\lambda}^{\left(\textrm{I}\Gamma\right)}\left(0\right)=0$ until it reaches an internal mode at $\tfrac{\lambda}{r+1}$. This greater flexibility, and therefore adaptability, of Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ is reflected in its versatility measure of $\mathcal{V}\left(F_{Y\mid r,\lambda}^{\left(\textrm{I}\Gamma\right)}\right)=3.1264$, which is equal to that of the $\textrm{Gamma}\left(r,\lambda\right)$ distribution and substantially greater than $\mathcal{V}\left(F_{Y\mid\alpha,\vartheta}^{\left(\textrm{P2}\right)}\right)=2.0874$.
[Uncaptioned image][Uncaptioned image]
Figure 2. PDFs of Pareto $\textrm{Pareto 2}\left(\alpha=1,\vartheta\right)$ and Inverse $\textrm{Inverse Gamma}\left(r=1,\lambda\right)$ Distributions (from left to right)
[variable parameter $=0.5$ for dotted line, $1.0$ for solid line, and $2.0$ for dashed line]
To investigate the relative effectiveness of the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ and Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ models in practice, we perform a series of four simple simulation experiments (I, II, III, and IV) to evaluate each model’s ability to distinguish itself from the other model within a hypothesis-testing framework:
  • In Experiment I, we take Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ as the null hypothesis, generate $10,000$ IID pairs $\left(\alpha_{i}\sim\textrm{Lognormal}\left(\mu=0,\sigma=1\right),\vartheta_{i}\sim\textrm{Lognormal}\left(\mu=0,\sigma=1\right)\right)$, and for each pair generate $n$ IID observations $Y_{j}\sim\textrm{Pareto 2}\left(\alpha_{i},\vartheta_{i}\right)$ (where $n\in\left\{50,100,500,1000,5000,\ldots,100,000\right\}$). We then use the $n$ IID $Y_{j}$ to perform a minimum chi-squared goodness-of-fit test (at the $0.05$ level of significance), [^23] In all minimum chi-squared tests, we employ maximum likelihood to estimate parameters and Sturges’ rule to set the number of test bins. and record the proportion of false rejections (out of the 10,000 $\left(\alpha_{i},\vartheta_{i}\right)$ pairs) as an estimate of Pr{Type 1 $\Pr\left\{\textrm{Type 1 Error}\right\}$.
  • In Experiment II, we take Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ as the null hypothesis, generate $10,000$ IID pairs $\left(r_{i}\sim\textrm{Lognormal}\left(\mu=0,\sigma=1\right),\lambda_{i}\sim\textrm{Lognormal}\left(\mu=0,\sigma=1\right)\right)$, and for each pair generate $n$ IID observations $Y_{j}\sim\textrm{Inverse Gamma}\left(r_{i},\lambda_{i}\right)$ (where $n\in\left\{50,100,500,1000,5000,\ldots,100,000\right\}$). We then use the $n$ IID $Y_{j}$ to perform a minimum chi-squared goodness-of-fit test (at the $0.05$ level), and record the proportion of false retentions (out of the 10,000 $\left(r_{i},\lambda_{i}\right)$ pairs) as an estimate of Pr{Type 2 $\Pr\left\{\textrm{Type 2 Error}\right\}$.
  • Experiments III and IV are analogous to Experiments I and II, respectively, with the roles of Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ and Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ reversed.
The experimental results in Table 5 show rather clearly that Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ performs better on the whole than Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$. Whereas both models yield satisfactorily low frequencies of Type 1 error (approaching the $0.05$ level of significance from below as $n$ increases), the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ hypothesis is consistently poor in recognizing and rejecting Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ observations (with Type 2 error rates above $0.80$). This shortcoming, which does not improve for larger sample sizes, appears to derive from the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$’s insufficient sensitivity to variations in the frequency of observations at the lower end of the sample space. In short, the model gives too much credence to observations consistent with the hypothesized tail behavior, thereby failing to discern the presence of otherwise non-Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ sample characteristics. Although too limited to be conclusive, the present analysis provides strong prima facie support for using the proposed versatility measure to identify and eschew the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ model’s idiosyncrasy.
Table 5. Hypothesis-Test Results from Simulated Data
I. $\boldsymbol{H_{0}}$: Pareto 2 II. $\boldsymbol{H_{0}}$: Pareto 2 III. $\boldsymbol{H_{0}}$: Inverse Gamma IV. $\boldsymbol{H_{0}}$: Inverse Gamma
(Pareto 2 Obs.) (Inverse Gamma Obs.) (Inverse Gamma Obs.) (Pareto 2 Obs.)
Sample Size (n) Pr{Type 1 Error} Pr{Type 2 Error} Pr{Type 1 Error} Pr{Type 2 Error}
$50$ $0.0000$ $1.0000$ $0.0034$ $1.0000$
$100$ $0.0056$ $0.8671$ $0.0131$ $0.9999$
$500$ $0.0181$ $0.8113$ $0.0195$ $0.6822$
$1000$ $0.0239$ $0.8084$ $0.0252$ $0.5129$
$5000$ $0.0296$ $0.8153$ $0.0362$ $0.3039$
$10,000$ $0.0320$ $0.8252$ $0.0379$ $0.2478$
$50,000$ $0.0402$ $0.8541$ $0.0435$ $0.1599$
$100,000$ $0.0482$ $0.8645$ $0.0494$ $0.1367$